There were 284 tickets sold for a basketball game. The activity cardholder's ticket cost $1.25 and the non cardholders' tickets cost $2.25. The total amount of money collected was $521. How many of each kind of ticket were sold?
How many activity cardholders ticket were sold?
How many non-cardholders ticket were sold?

Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

let x be the number of activity cardholder's tickets sold at $1.25
non - 284-x tickets $2.25
Total revene = $521
..
1,25x+(284-x)*2.25 = 521
1.25x + 639-2.25x=521
-x= -639-521
x= 118 activity cardholder's tickets
284-118= 166 non activity cardholder's tickets

Answer by rfer(16322)

c+n=284 c=284-n
1.25c+2.25n=521
1.25(284-n)+2.25n=521
355-1.25n+2.25n=521
n=166
c=284-166
c=118
-----------------
1.25*118+2.25*166=521
147.50+373.50=521